find the mass of carbon that contain the same number of atoms as contained by 560g of fe
 

Grade:11

1 Answers

Shivam jha
6 years ago
Number of moles of C =560/56=10mol So number of atom=10×6.022×10 power 23. So 12g of C contain 1 mole therfore fo 10mol 120g is required for same atom.